Two point charge placed at a distance of 20 cm in ai, repel each other with a certain force. When a dielectric slab of thickness 8 cm and dielectric constant K is introduced between these point charges, the force becomes ½ of its previous value. Then K is approximately equal to? 2 4 2 ½ 1

Two point charge placed at a distance of 20 cm in ai, repel each other with a certain force. When a dielectric slab of thickness 8 cm and dielectric constant K is introduced between these point charges, the force becomes ½ of its previous value. Then is approximately equal to?
  1. 2
  2. 4
  3. 2½
  4. 1

Grade:12

1 Answers

Arun
4 years ago
  • We know F=1/4pi£0 q1q2/r^2
  • F/2=1/4pi£0 q1q2/(distance bw charges without dielectric+thickness of slab under root of dielectric constant-1)^2
  • Substituting nd solving gives the value 4